OpenStudy (anonymous):

If the roots of the equation a(x)^2 + bx + c = 0, are in the ratio m:n, prove that: mn . (b)^2 = ac (m+n)^2

OpenStudy (anonymous):

I did it in the following way. Is it correct?

OpenStudy (anonymous):

Let p,q be the roots. We know that, p + q = -(b/a) And, pq = c/a Given , p/q = m/n So, LHS = mn. b. b = b^2 . c/a Now, RHS = ac (m+n)(m+n) =ac. [(-b)/a]^2 = b^2.c/a So, LHS = RHS proved. Is it correct?

OpenStudy (kinggeorge):

That looks correct to me.

OpenStudy (anonymous):

Really? I want to get it confirmed from others also.

OpenStudy (kinggeorge):

The only thing that wasn't immediately clear to me, was why mn=c/a, and why m+n=(-b/a). If you explained that more clearly, I would be an excellent proof.

OpenStudy (anonymous):

See, I have taken, p/q = m/n We know that, in a quadratic equation, sum of the roots = -b/a Product of roots = c/a

OpenStudy (kinggeorge):

It's the substitutions you were doing for m and n that I didn't immediately see. Like I said, if you elaborated on that part of the proof, it would be an excellent proof.

OpenStudy (anonymous):

Okay. Vishal Bhai, is it correct? Its a 4 mark question.

OpenStudy (vishal_kothari):

let me see it properly...

OpenStudy (anonymous):

Yeah.

OpenStudy (vishal_kothari):

oh no light gayi....but i can tell use the quadratic formula....

OpenStudy (anonymous):

I didn't understand.

OpenStudy (vishal_kothari):

truly speaking india needs development....

OpenStudy (vishal_kothari):